OpenStudy (anonymous):

Find the inverse of y=4x=1/3x-2. state the domain and range of the function and its inverse.Please show work

OpenStudy (anonymous):

sorry it is y=4x+1/3x-2

OpenStudy (anonymous):

you have 2 equal sigh=ns

OpenStudy (anonymous):

ok so x = 4y + 1/3 y - 2 is the inverse now sollve it for y

OpenStudy (anonymous):

since the original function is linear its domain and range are x \[\epsilon\]R and also y ... so in the invewrse the domain and range remain the same

OpenStudy (zarkon):

is the function \[y=\frac{4x+1}{3x-2}\]?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

then you should write it as y=(4x+1)/(3x-2) so there is confusion.

OpenStudy (zarkon):

*is no confusion

OpenStudy (anonymous):

yes confusion

OpenStudy (zarkon):

writing 4x+1/3x-2 is the same as \[4x+\frac{1}{3}x-2\]

OpenStudy (anonymous):

my book just states find the invers of y=4x+1/3x-2. state the domain and range of the function and its inverse

OpenStudy (anonymous):

no it is divided by

OpenStudy (anonymous):

does that make sence

OpenStudy (anonymous):

it is stated y= 4x+1 divided by 3x-2

OpenStudy (zarkon):

ic

OpenStudy (zarkon):

solve \[x=\frac{4y+1}{3y-2}\]for \[y\]

OpenStudy (anonymous):

yes and I am not sure how. I am so new at this I am confused

OpenStudy (zarkon):

\[x=\frac{4y+1}{3y-2}\] \[x(3y-2)=4y+1\] \[3xy-2x=4y+1\] \[3xy-4y=2x+1\] \[(3x-4)y=2x+1\] \[y=\frac{2x+1}{3x-4}\]

OpenStudy (anonymous):

Thank you now I have one to look at to do the rest.